# Instruction-based Time Series Editing

### v3

@@ Abstract.
In time series editing, we aim to modify some properties of a given time series without altering others. For example, when analyzing a hospital patient’s blood pressure, we may add a sudden early drop and observe how it impacts their future, while preserving other conditions. Existing diffusion-based editors rely on rigid, predefined attribute vectors as conditions and produce all-or-nothing edits through sampling. This attribute- and sampling-based approach limits flexibility in condition format and lacks customizable control over editing strength. To overcome these limitations, we introduce Instruction-based Time Series Editing, where users specify intended edits using natural language. This allows users to express a wider range of edits in a more accessible format. We then introduce InstructTime, the first instruction-based time series editor. InstructTime takes in time series and instructions, embeds them into a shared multi-modal representation space, then decodes their embeddings to generate edited time series. By learning a structured multi-modal representation space, we can easily interpolate between embeddings to achieve varying degrees of edit. To handle local and global edits together, we propose multi-resolution encoders. In our experiments, we use synthetic and real datasets and find that InstructTime is a state-of-the-art time series editor: InstructTime achieves high-quality edits with controllable strength, can generalize to unseen instructions, and can be easily adapted to unseen conditions through few-shot learning. All of our code is publicly-available: https://github.com/JiaxingQiu/InstructTime.
Time Series Editing, Natural Language Instruction-based Editing, Shared Embedding Space, Contrastive Learning.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM SIGKDD; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Modeling methodologiesccs: Computing methodologies Learning latent representations
@@ 1. Introduction
State-of-the-art TSE methods focus on attribute-level control and use conditional diffusion (Jing et al., 2024; Narasimhan et al., 2024). However, these methods have two limitations. Firstly, they represent attributes as continuous or categorical feature vectors. These attributes require handcrafted feature engineering, constrains all future edits to capture only these attributes, and limits users to this hardcoded interaction with the editor. Secondly, diffusion-based methods perform all-or-nothing edits by sampling diverse outputs, offering no control over editing strength. In contrast, allowing edits with varying condition influence enables finer user control and supports insight and hypothesis generation—e.g., gradually adding abnormality to a normal heart rate from mild to severe.
@@ 3.2. InstructTime
To efficiently compress an instruction into a semantic embedding, we firstly leverage a pretrained text embedding model (frozen parameters) to encode the instruction $\mathbf{c}$ into a numeric vector. Such vector captures rich semantic meaning and paraphrastic similarity of the multiple conditions described in the instruction (Reimers and Gurevych, 2019; Jiang et al., 2025). The vector is then processed by $k$ parallel multi-layer perceptrons (MLPs), each predicting a $d$-dimensional vector $\mathbf{z}_{c_{k}}$. These vectors are concatenated to form the final $D$-dimensional instruction embedding, $\mathbf{z}_{c}=[\mathbf{z}_{c_{1}}\,\|\,\mathbf{z}_{c_{2}}\,\|\,\cdots\,\|\,\mathbf{z}_{c_{k}}]$, where $D=k\cdot d$. This chunked representation enables resolution-wise semantic alignment—where the $r^{\text{th}}$ segment of the instruction embedding aligns with the $r^{\text{th}}$ segment of the time series embedding ($r={1,\ldots,k}$)—thereby linking each segment’s semantic representation to its corresponding temporal resolution during both contrastive fusion and decoding. A detailed explanation of how the contrastive learning encourages resolution-wise semantic alignment is provided in Appendix B. Lastly, $\mathbf{z}_{c}$ is also normalized to have unit length to ensure it resides in the same bounded embedding space as $\mathbf{z}_{x}$.
@@
$\mathcal{L}_{\text{recon}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathrm{x}_{i}-\hat{\mathrm{x}}_{i}\right\|_{2}^{2}.$
@@ 4.1. Experimental Setup
Instruction-based
Time Weaver -11.05 5.93 0.63 3.50 1.54 -0.97 0.17 0.66 -0.15 0.13 1.03
@@
Attribute-based
Time Weaver -12.19 6.65 0.15 3.58 1.54 -1.26 0.31 0.73 -1.55 0.15 1.40
@@
where $p(\tilde{\mathrm{c}}|\mathrm{x})$ is the probability of $\hat{\mathrm{x}}$ being classified to the correct level of attribute $\tilde{\mathrm{c}}$. For each attribute, we compute $p(\tilde{\mathrm{c}}|\mathrm{x})$ using a multi-class classifier that appends a softmax prediction head to the architecture of time series encoder $\mathcal{E}_{\phi}$, trained end-to-end on the held-out dataset. Higher $\text{RaTS}(\hat{\mathrm{x}},\mathrm{x},\tilde{\mathrm{c}}_{\text{edit}})$ indicates better editability. The absolute value $|\text{RaTS}(\hat{\mathrm{x}},\mathrm{x},\tilde{\mathrm{c}}_{\text{prsv}})|$ measures how well the output $\hat{\mathrm{x}}$ preserves the remaining attributes $\tilde{\mathrm{c}}_{\text{prsv}}$ of the input $\mathrm{x}$; values closer to zero indicate better preservability (Jing et al., 2024). Note that we use an end-to-end trained classifier to calculate $p(\tilde{\mathrm{c}}|\mathrm{x})$ for each attribute, rather than using the retrieval softmax probability computed from the shared embedding space as done by (Jing et al., 2024), because our generation process used the shared embedding space, and using retrieval softmax probability may favor edits by our method.
@@ 4.3. Interpolated Editing
Instruction-based
(unseen-phrasing)
Time Weaver -12.95 5.24 0.57 3.25 1.48 -0.97 0.14 0.47 -0.27 0.15 1.34
@@
As the editing strength $w$ increases, the edited time series should progressively reflect the target edited attributes $\tilde{\mathbf{c}}_{\mathrm{edit}}$ to a greater extent, while maintaining high preservability of $\tilde{\mathbf{c}}_{\mathrm{prsv}}$. Figure 3 shows a case study comparing the editing procedures by InstructTime and TEdit. As editing strength $w$ increases, InstructTime gradually turns a flat trend into upward trends with increasing slope, while preserving local shifts or global seasonality. In contrast, TEdit produces upward trends with random slopes and scattered distances from the original while preserving other attributes. Figure 4 shows the progressively change in editability ($\small\Delta\text{DTW}$, RaTS) and preservability ($\small|\text{RaTS}|$) by InstructTime as the editing strength $w$ increases, compared to the fixed performances by TEdit and Time Weaver. As $w$ increases, the target attributes become more prominent in the edited time series, resulting in stronger editability. Larger $w$ also leads to greater divergence from the input time series and lower preservability, as changes to target attributes may introduce turbulence in other attributes meant to be preserved. Despite this inherent trade-off, InstructTime maintains comparable preservability as $w$ approaches 1 and achieves better preservability in the low-$w$ regime than other methods. Compared to attribute-based editing, InstructTime maintains high editing quality in the instruction-based setting, while other methods show reduced editability and preservability, as indicated by higher $\small\Delta\text{DTW}$, lower RaTS, and higher $\small|\text{RaTS}|$.
@@ 4.4. Generalizability to Unseen Instructions
Figure 6 visualizes the re-calibration of time series and instruction embeddings in the shared embedding space after after tuning with a single example (one-shot). Before tuning (left panel), the model is trained on two trend levels: flat (red) and downward (blue). Dots represent embeddings of 100 time series per level; triangles denote their instruction embeddings. The instruction and time series embeddings align well for the seen levels, but for the unseen “upward trend” (green), the triangle and dots are far apart in the embedding space. The model is then tuned on a small set of instruction–time series pairs, generated by providing an example pair from the unseen level “upward trend” and editing it toward the seen levels to create a few synthetic samples for tuning. This allows the model to adjust its internal representation of the unseen condition in terms of its relative position and distance to the seen conditions in the embedding space. As shown in the right panel, the time series and instruction embeddings are well calibrated across all three levels, highlighting InstructTime’s ability to generalize under limited supervision.
@@ 6. Acknowledgements
We thank University of Virginia’s High Performance Computing team for computing resources. We thank the reviewers for very constructive feedback. TH was supported in part by a CapitalOne faculty fellowship and the University of Virginia’s National Security Data & Policy Institute through the U.S. Department of Defense Contracting Activity #2024-24070100001.

### v4

@@ Abstract.
In time series editing, we aim to modify some properties of a given time series without altering others. For example, when analyzing a hospital patient’s blood pressure, we may add a sudden early drop and observe how it impacts their future, while preserving other conditions. Existing diffusion-based editors rely on rigid, predefined attribute vectors as conditions and produce all-or-nothing edits through sampling. This attribute- and sampling-based approach limits flexibility in condition format and lacks customizable control over editing strength. To overcome these limitations, we introduce Instruction-based Time Series Editing, where users specify intended edits using natural language. This allows users to express a wider range of edits in a more accessible format. We then introduce InstructTime, the first instruction-based time series editor. InstructTime takes in time series and instructions, embeds them into a shared multi-modal representation space, then decodes their embeddings to generate edited time series. By learning a structured multi-modal representation space, we can easily interpolate between embeddings to achieve varying degrees of edit. To handle local and global edits together, we propose multi-resolution encoders. In our experiments, we use synthetic and real datasets and find that InstructTime is a state-of-the-art time series editor: InstructTime achieves high-quality edits with controllable strength, can generalize to unseen instructions, and can be easily adapted to unseen conditions through few-shot learning.
Time Series Editing; Natural Language Instruction; Shared Embedding Space; Contrastive Learning
submissionid: v1rtp0903journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780299isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Modeling methodologiesccs: Computing methodologies Learning latent representations
Resource Availability:
The source code of this paper has been made publicly available at https://github.com/JiaxingQiu/InstructTime.
@@ 1. Introduction
State-of-the-art TSE methods focus on attribute-level control and use conditional diffusion (Jing et al., 2024; Narasimhan et al., 2024). However, these methods have two limitations. Firstly, they represent attributes as continuous or categorical feature vectors. These attributes require handcrafted feature engineering, and constrain all future edits to capture only these attributes, and limits users to this hardcoded interaction with the editor. Secondly, diffusion-based methods perform all-or-nothing edits by sampling diverse outputs, offering no control over editing strength. In contrast, allowing edits with varying condition influence enables finer user control and supports insight and hypothesis generation—e.g., gradually adding abnormality to a normal heart rate from mild to severe.
@@ 3.2. InstructTime
To efficiently compress an instruction into a semantic embedding, we first leverage a pretrained text embedding model (frozen parameters) to encode the instruction $\mathbf{c}$ into a numeric vector. Such vector captures rich semantic meaning and paraphrastic similarity of the multiple conditions described in the instruction (Reimers and Gurevych, 2019; Jiang et al., 2025). The vector is then processed by $k$ parallel multi-layer perceptrons (MLPs), each predicting a $d$-dimensional vector $\mathbf{z}_{c_{k}}$. These vectors are concatenated to form the final $D$-dimensional instruction embedding, $\mathbf{z}_{c}=[\mathbf{z}_{c_{1}}\,\|\,\mathbf{z}_{c_{2}}\,\|\,\cdots\,\|\,\mathbf{z}_{c_{k}}]$, where $D=k\cdot d$. This chunked representation enables resolution-wise semantic alignment—where the $r^{\text{th}}$ segment of the instruction embedding aligns with the $r^{\text{th}}$ segment of the time series embedding ($r={1,\ldots,k}$)—thereby linking each segment’s semantic representation to its corresponding temporal resolution during both contrastive fusion and decoding. A detailed explanation of how the contrastive learning encourages resolution-wise semantic alignment is provided in Appendix B. Lastly, $\mathbf{z}_{c}$ is also normalized to have unit length to ensure it resides in the same bounded embedding space as $\mathbf{z}_{x}$.
@@
$\small\mathcal{L}_{\text{recon}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathrm{x}_{i}-\hat{\mathrm{x}}_{i}\right\|_{2}^{2}.$
@@ 4.1. Experimental Setup
Instruction-based Time Weaver -11.05 5.93 0.63 3.50 1.54 -0.97 0.17 0.66 -0.15 0.13 1.03
@@
Attribute-based Time Weaver -12.19 6.65 0.15 3.58 1.54 -1.26 0.31 0.73 -1.55 0.15 1.40
@@
where $p(\tilde{\mathrm{c}}|\mathrm{x})$ is the probability of $\hat{\mathrm{x}}$ being classified to the correct level of attribute $\tilde{\mathrm{c}}$. For each attribute, we compute $p(\tilde{\mathrm{c}}|\mathrm{x})$ using a multi-class classifier that appends a softmax prediction head to the architecture of time series encoder $\mathcal{E}_{\phi}$, trained end-to-end on the held-out dataset. Higher $\text{RaTS}(\hat{\mathrm{x}},\mathrm{x},\tilde{\mathrm{c}}_{\text{edit}})$ indicates better editability. The absolute value $|\text{RaTS}(\hat{\mathrm{x}},\mathrm{x},\tilde{\mathrm{c}}_{\text{prsv}})|$ measures how well the output $\hat{\mathrm{x}}$ preserves the remaining attributes $\tilde{\mathrm{c}}_{\text{prsv}}$ of the input $\mathrm{x}$; values closer to zero indicate better preservability (Jing et al., 2024). Note that we use an end-to-end trained classifier to calculate $p(\tilde{\mathrm{c}}|\mathrm{x})$ for each attribute, rather than using the retrieval softmax probability computed from the shared embedding space as done by (Jing et al., 2024), because our generation process used the shared embedding space, and using retrieval softmax probability may favor our edits. Our results remain consistent across runs, while uncertainty in metrics can occur due to the sampling procedure in prior methods.
@@ 4.3. Interpolated Editing
Instruction-based (unseen-phrasing) Time Weaver -12.95 5.24 0.57 3.25 1.48 -0.97 0.14 0.47 -0.27 0.15 1.34
@@
As the editing strength $w$ increases, the edited time series should progressively reflect the target edited attributes $\tilde{\mathbf{c}}_{\mathrm{edit}}$ to a greater extent, while maintaining high preservability of $\tilde{\mathbf{c}}_{\mathrm{prsv}}$. Figure 3 shows a case study comparing the editing procedures by InstructTime and TEdit. As editing strength $w$ increases, InstructTime gradually turns a flat trend into upward trends with increasing slope, while preserving local shifts or global seasonality. In contrast, TEdit produces upward trends with random slopes and scattered distances from the original while preserving other attributes. Figure 4 shows the progressive change in editability ($\small\Delta\text{DTW}$, RaTS) and preservability ($\small|\text{RaTS}|$) by InstructTime as the editing strength $w$ increases, compared to the fixed performances by TEdit and Time Weaver. As $w$ increases, the target attributes become more prominent in the edited time series, resulting in stronger editability. Larger $w$ also leads to greater divergence from the input time series and lower preservability, as changes to target attributes may introduce turbulence in other attributes meant to be preserved. Despite this inherent trade-off, InstructTime maintains comparable preservability as $w$ approaches 1 and achieves better preservability in the low-$w$ regime than other methods. Compared to attribute-based editing, InstructTime maintains high editing quality in the instruction-based setting, while other methods show reduced editability and preservability, as indicated by higher $\small\Delta\text{DTW}$, lower RaTS, and higher $\small|\text{RaTS}|$.
@@ 4.4. Generalizability to Unseen Instructions
Figure 6 visualizes the re-calibration of time series and instruction embeddings in the shared embedding space after tuning with a single example (one-shot). Before tuning (left panel), the model is trained on two trend levels: flat (red) and downward (blue). Dots represent embeddings of 100 time series per level; triangles denote their instruction embeddings. The instruction and time series embeddings align well for the seen levels, but for the unseen “upward trend” (green), the triangle and dots are far apart in the embedding space. The model is then tuned on a small set of instruction–time series pairs, generated by providing an example pair from the unseen level “upward trend” and editing it toward the seen levels to create a few synthetic samples for tuning. This allows the model to adjust its internal representation of the unseen condition in terms of its relative position and distance to the seen conditions in the embedding space. As shown in the right panel, the time series and instruction embeddings are well calibrated across all three levels, highlighting InstructTime’s ability to generalize under limited supervision.
@@ 6. Acknowledgements
We thank University of Virginia’s High Performance Computing team for computing resources. We thank the reviewers for their constructive feedback, which helps improve the manuscript in many important ways. Thomas Hartvigsen was supported in part by a CapitalOne faculty fellowship and the University of Virginia’s National Security Data and Policy Institute through the U.S. Department of Defense Contracting Activity, Award Number #2024-24070100001. Contact author: Jiaxing Qiu (jq2uw@virginia.edu).